The Optical and Optical/Stark Spectrum of Iridium Monocarbide and Mononitride

A. J. Marr, M. E. Flores and \underline{T. C. Steimle}, Dept. of Chemistry and Biochemistry, Arizona State University, Tempe, AZ 85287-1604.

Supersonic molecular beam samples of iridium monocarbide, IrC, and iridium mononitride, IrN, were generated using a laser ablation/reaction source and characterized using high resolution ( Δν < 30 MHz FWHM) laser induced fluorescence spectroscopy. This is the first identification of gaseous IrN. The conventional high temperature spectra of IrC was recorded some time ago^{a}^{,}^{b}. Numerous strong band systems in the 18800 cm^{−1} to 14360 cm^{−1} spectral range were assigned as the (v′,0) progression of the Ã^{1}Π - X̃^{1}Σ^{+} band system of IrN. The (1,0) and (0,0) bands were analyzed to produce a set of fine and hyperfine parameters. The electric field induced effects on the R(0) branch features of these bands were analyzed to produce permanent electric dipole moments: Ã^{1}Π(v=0) 2.78(2) D, Ã^{1}Π(v=1) 2.64(2) D and X̃^{1}Σ^{+}(v=0) 1.66(1) D. The (0,0) D̃^{2}Φ_{7/2}-X̃^{2}Δ_{5/2} band system of IrC was recorded and analyzed to produce a set of fine and hyperfine parameters. The electric field induced effects on the R(2.5) branch feature were analyzed to produce permanent electric dipole moments: D̃^{2}Φ_{7/2} (v=0) 2.61(6) D and X̃^{2}Δ_{5/2} (v=0) 1.60(7) D. Plausible electronic configurations consistent with the experimental observations are given.

aK. Jansson, R. Scullman and B. Yttermo, Chem. Phys. Lett. 4, 188 (1969).

bK. Jansson and R. Scullman, J. Mol. Spectrosc. 36, 248 (1970).

Time required: 15 min Session in which paper is recomended for presentation: 8